OpenStudy (anonymous):

Marta’s checkbook shows that she wrote a check for $987 for building materials. What was the price of the materials if the sales tax is 5%? @Pre-Algebra

OpenStudy (anonymous):

show ur work plz

OpenStudy (anonymous):

987 + 5/100 * 987

OpenStudy (anonymous):

$ 940

OpenStudy (anonymous):

what huh???

OpenStudy (anonymous):

x + (5/100)x = 987 (21x)/20 = 987 x = 19740/21 x = 940 $

OpenStudy (anonymous):

i dnt get it???

OpenStudy (anonymous):

Marta bought those material's for $987. That is cost of those building material + sales tax(5%) Let cost of building material's be x. Then x + 5% of x = 987 x + (5/100) x =987 x + (x/20) = 987 (21x)/20 = 987 21x = 987*20 = 19740 x = 19740/21 x = 940 Cost of building material = $ 940

OpenStudy (anonymous):

You can check it. 940 + (5/100)*940 = 987

OpenStudy (anonymous):

if he is paying $987 dollars for building material then you would change your precent to a decimal that becomes .05 then you multiply it together 987x.05 which equals 49.35 thats your sales tax then you add that on too 987 ,, 987+49.35 which equals 1036.35.